OpenStudy (anonymous):

Use properties of logarithms to expand the logarithmic expression as much as possible. Where possible, evaluate logarithmic expressions without using a calculator. log (1000000x)=

terenzreignz (terenzreignz):

This is log base 10, right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Use this property \[\huge \log(MN) = \log(M) + \log(N)\]

OpenStudy (anonymous):

wouldn't that be log(1000000)+log(x) because if so that's what I thought originally and it was wrong

terenzreignz (terenzreignz):

Yes, it would. Now you can simplify one of these logs...

OpenStudy (anonymous):

i know it would be the log(1000000) just not sure what it would simplify to